OpenStudy (anonymous):

A = 2(lw + wh + lh), for l given, w = 5, h = 3, and A = 158

OpenStudy (anonymous):

ok so plug in your given values A = 2(lw + wh + lh) 158=2(l(5)+5(3)+l(3)) simplify 158=2(5l+15+3l) get rid of the parenthesis by multiplying everything in them by 2 158=10l+30+6l combine like terms 158=16l+30 128=16l divide both sides by 16 l=8

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

No problem anytime:)

OpenStudy (anonymous):

Hello